### 02 Dec 2023 v1

# Industrial Internet of Things Intelligence Empowering Smart Manufacturing: A Literature Review

Abstract
The fiercely competitive business environment and increasingly personalized customization needs are driving the digital transformation and upgrading of the manufacturing industry. IIoT intelligence, which can provide innovative and efficient solutions for various aspects of the manufacturing value chain, illuminates the path of transformation for the manufacturing industry. It’s time to provide a systematic vision of IIoT intelligence. However, existing surveys often focus on specific areas of IIoT intelligence, leading researchers and readers to have biases in their understanding of IIoT intelligence, that is, believing that research in one direction is the most important for the development of IIoT intelligence, while ignoring contributions from other directions. Therefore, this paper provides a comprehensive overview of IIoT intelligence. We first conduct an in-depth analysis of the inevitability of manufacturing transformation and study the successful experiences from the practices of Chinese enterprises. Then we give our definition of IIoT intelligence and demonstrate the value of IIoT intelligence for industries in fucntions, operations, deployments, and application. Afterwards, we propose a hierarchical development architecture for IIoT intelligence, which consists of five layers. The practical values of technical upgrades at each layer are illustrated by a close look on lighthouse factories. Following that, we identify seven kinds of technologies that accelerate the transformation of manufacturing, and clarify their contributions. Finally, we explore the open challenges and development trends from four aspects to inspire future researches.
Index Terms:
Industrial Internet of Things, IIoT intelligence, Smart manufacturing, Artificial intelligence
I Introduction
Manufacturing holds an important position in the global economy. It accounts for 16% of global GDP, and its research and development (R&D) spend accounts for 64% of global R&D expenditure [1]. However, manufacturing is suffering from many troubles. The stagnant productivity, increasingly personalized customization requirements, long overdue innovation, rapidly shrinking humanpower, and expectation for one-stop comfortable service, all of these are pushing manufacturing to upgrade and transform. At the same time, the rapid development of intelligent technologies in recent year also inspires governments and enterprises to carry out digitization transformation. Under the circumstance, governments issued a series of policies [2, 3, 4, 5], such as New Robot Strategy, Industrial Strategy 2030, etc. Inspired by the policy dividend, enterprises begin to take the path of industrial revolution and innovation [6][7].
Industrial Internet of Things (IIoT), which makes all types of industrial equipment to be connected together through networks and can further support data collection, exchange and analysis [8], have realized great achievements for improve the digitalization level of industries. To make the research values of IIoT clear, some works surveyed the various definitions of IIoT and enabling technologies, and proposed open challenges to inspire future research directions. Boyes et al. [9] differentiated the connotation of IIoT with industry 4.0, cyber physical system, industrial automation & control systems, supervisory constrol and data acquisition, industrial internet, and emphasized interconnection among smart objects, cyber-physical assets and computing platforms. Khan et al. [10] reviewed the research efforts in the areas of architectures and frameworks, communication protocols and data management schemes for IIoT. The general architecture for IIoT system in [10] includes networked entities, cloud/edge computing platforms, and enterprise business domains. The researches of IIoT communication protocols are highlighted. Cloud/Edge computing is used to provide distributed high performance computing resources for IIoT. Malik et al. [11] gave the architecture of IIoT, which consists of physical layer, data link layer, network and Internet layer, transport layer and application layer, and analyzed the potential applications of IIoT, including environment monitoring, smart agriculture, healthcare, supply chain management and home automation. The common insight of these works for future researches is to combine more advanced technologies with IIoT, such as blockchain, big data analysis, deep learning, etc., so that IIoT can become more intelligent.
As expected, in recent years, the integration of IIoT with many other advanced technologies makes IIoT emerge intelligence in terms of ubiquitous reliable communication, dynamic environment adaptation, flexible manufacturing, etc. This also accelerates the transformation progress of manufacturing industries. To be specific, when IIoT is combined with 5G, its demands of multiple device connectivity and high data rate, more bandwidth, low-latency quality of service (QoS) can be easily satisfied. Chettri et al. [12] and Mahmood et al. [13] provided detailed overview of challenges and vision of various communication industries in 5G IIoT systems. Time sensitive networking (TSN) is a set of IEEE 802.1 standards that defines mechanisms to provide deterministic services through IEEE 802 networks [14, 15, 16]. Deterministic services include guaranteed packet transport with bounded latency, low packet delay variation, and low packet loss. Combined with TSN, IIoT can ensure the deterministic response for network traffics between any two devices. Bello et al. [17] described TSN protocols in detail and provided some cases while using TSN into IIoT. Deep learning has potentials to be combined with IIoT as well. Khalil et al. [18] presented the various deep learning techniques, including convolutional neural networks, autoencoders, and recurrent neural networks, as well as their use in different industries. The product R&D cost in IIoT can be reduced through introducing digital experience [19][20]. The labour efficiency grows with the help of remote monitoring and control [21]. The industrial robots can gradually replace human in some physically demanding and routine jobs [22][23].
It’s time to have a systematic understanding for IIoT intelligence. However, existing surveys always focused on IIoT intelligence in specific domain, such as 5G in IIoT [12, 13], TSN in IIoT [17], deep learning in IIoT [18, 24, 25], edge computing in IIoT [8, 26, 27]. This will lead to biases among researchers/readers in their understanding of IIoT intelligence, that is, believing that the research in one direction is the most important for the development of IIoT intelligence, while neglecting the contributions of other directions. Under the circumstance, the contributions/positions of advanced technologies in the architecture of IIoT intelligence cannot be clarified, either.
Therefore, in this paper, we provide a comprehensive overview of IIoT intelligence. Considering that the development of IIoT intelligence integrates the wisdom of governments, enterprises, and researchers, we expand the information sources from academic papers to government documents, industry reports, and other related literature. The essential of this paper is to answer four questions as follows.
  • What is IIoT intelligence and what can IIoT intelligence do for smart manufacturing?
  • What is the development architecture of IIoT intelligence?
  • Which technologies are accelerating the development of IIoT intelligence?
  • What are the open challenges and future trends of IIoT intelligence?
In response to the first question, we conduct an in-depth analysis of the inevitability of manufacturing transformation under the pressure of customization demand, increased labor costs, and competitive environment, as well as the incentives of government policies and rapid development of advanced technology. Then, based on the practical exploration of Chinese industries, the secrets of successful transformation are revealed, including building an effective production architecture, building a customer-centric value chain, building innovation business models, building digital capabilities of manufacturing and organization. The secrets inspire the emergence and development of IIoT intelligence. We then give our definition of IIoT intelligence, i.e. a series of techniques, methods, productions and platforms that can been taken within the entire value chain of manufacturing to build autonomous capabilities. The value chain of manufacturing involves R&D, production, operation and maintenance, marketing, management, services, etc. The built autonomous capabilities through IIoT intelligence include digital connection and perception, intelligent analysis and cognition, immediate decisions, etc. Combining with practice, we demonstrate the value of IIoT intelligence for manufacturing industries in functions, operations, deployments and application scenarios.
To answer the second question, we propose a hierarchical development architecture to build a system-level understanding about how IIoT intelligence empowers smart manufacturing. Specifically, the architecture consists of five layers. Each layer has own missions. The equipment layer builds the basis of automatic production. The network layer connects human-machine-things, and overcomes isolated islands of information. The software layer provides digital representation of industrial processes. The modeling layer makes digital twin for physical processes to connect virtual and physical spaces. The analysis and optimization layer is responsible for mining industrial big data and optimizing industrial processes. Each layer of IIoT intelligence has its own key technologies, making its capabilities stronger. Interactions will occur between levels to fully enhance the intelligence and digitization level of the manufacturing industry. In addition, we also closely observe the lighthouse factory [28][29] to demonstrate that technological upgrades at each layer enhance the company’s returns.
To deal with the third question, we identify seven kinds of technologies and clarify their contributions in accelerating the development of IIoT intelligence. To be specific, (a) industrial robots can become intelligent labors; (b) machine vision systems have the potential to be the eyes of industries; (c) networking facilitates ubiquitous connectivity; (d) digital twins connects cyber spaces and physical spaces; (e) deep learning boosts intelligence; (f) smart hardware builds foundation of intelligence; (g) cloud/edge computing inspires novel business and deployment.
To reply to the fourth question, we explore open challenges and future trends in four directions, based on the vision of the future manufacturing industry, including digital control, deterministic response, cost-friendly operation and deployment, IIoT intelligence proliferation.
The main contributions of this paper can be summarized as follows.
  • This paper provides a comprehensive literature review on IIoT intelligence to establish a systematic understanding of the positive impact of IIoT intelligence on smart manufacturing.
  • This paper proposes a hierarchical development architecture for IIoT intelligence and combines it with practical lighthouse factories to demonstrate the research value of various levels in improving enterprise returns.
  • This paper identify seven kinds of non-trivial technologies and clarify their contributions in accelerating the development of IIoT intelligence.
  • This paper discovers open challenges and predicts future trends of IIoT intelligence to inspire future researches.
This paper is structured as follows. Section II presents background and transformation secrets of smart manufacturing. The connotation of IIoT intelligence is explained in Section III. Section IV provides a reference hierarchical development architecture to build a systematic understanding for IIoT intelligence. Section V demonstrates the significant effects of IIoT intelligence through case studies on lighthouses and gradually released collaborations. Section VI identities some of the most influential technologies that facilitate IIoT intelligence. Section VII summarizes open challenges and future trends of IIoT intelligence. The conclusion is given in Section VIII.
II In-depth Perspectives on Smart Manufacturing
In recent years, the Forth Industrial Revolution has attracted attentions of governments, enterprises and universities, and has been fully launched. Many manufacturers are exploring the way to intelligent transformation.
II-A Background
In this section, we will answer one important question, why the manufacturing industry needs to undergo intelligent transformation, based on the factors in social environment and technical development.
In recent years, the population of many countries increases slowly, and meanwhile the workforce costs of manufacturing industries are rising. For example, according to the results of the Chinese National Bureau of Statistics[30] in Fig. 1, the workforce in manufacturing has been falling from 2014 to 2019, while workforce costs have been rising significantly. In addition, more and more customers are pursuing personalized customization and patient experience, which makes the architecture of manufacturing systems become more complex at every phase of end-to-end product life cycle throughout designing, production, and after service, as shown in Fig. 2. The capabilities of mass-producing customized products on demand and reacting rapidly to demand fluctuation will provide the essential competitive for companies. However, there are some challenges to build the capabilities, including adjusting the organizational structure, operating, and business modes. In general, the rising workforce costs and customization demands force the manufacturing community to embark intelligent transformation, so as to increase productivity and realize appealing returns.
Figure 1: The workforce in Chinese manufacturing has been falling from 2014 to 2019, while workforce costs have been rising significantly.
Figure 2: Personalized customization and patient experience enhance complexity of the architectures for manufacturing systems.
Recent years have witnessed the rapid development of a large number of advanced intelligent technologies, including cloud computing, 5G connection, industrial data analysis, industrial internet of things, etc. These technologies light the way for upgrading manufacturing. Many countries around the world also support intelligent transformation of manufacturing in the perspectives of policies. Around 20 economies released approximately 100 strategic planning or policy documents on artificial intelligence in the past three years by 2019, among which more than 50 documents are related with industrial application. For example, the governments of the United States, Japan, Germany, China have respectively launched The National Artificial Intelligence Research and Development strategic Plan [2], New Robot Strategy [3], Industrial Strategy 2030: Guidelines for a German and European industrial policy [4], Made in China 2025 [5]. In the government plans, some research domains are highlighted, such as product life cycle optimization, advanced robot, automatic driving, big data mining, cloud computing and other industrial applications.
II-B Secrets to smart manufacturing
In this section, we will answer the question, how to accelerate the progress of intelligent transformation for manufacturing industry, based on the practical experiences of Foxconn Industrial Internet, which is a famous enterprise that has realized a high-level intelligent transformation in China.
Foxconn Industrial Internet summarizes its recipes into three pillars and two capabilities [31], as presented in Fig. 3.
  • Building an effective production architecture to achieve intelligent transformation. Specifically, companies should improve flexible automated production lines, connect production data, and promote intelligent applications. As a result of these efforts, companies will improve their yield and throughput, energy efficiency and product quality, and reduce product cost, inventory, operating cost.
  • Building a customer-centric value chain. Companies are expected to develop new products and provide supporting service on the basis of accurately understanding customer needs. Enterprises may focus on optimizing different value chains, such as the value chains from suppliers to marketing, from enterprise management to after-sales service, from product development to consumers, etc.
  • Building innovation business models to enable digital manufacturing ecosystem through standardized collaboration models with academia, industrial partners and government, such as technology ecosystem [32], investment binding.
  • Building digital capabilities of manufacturing by combining traditional and new technologies. Although these technologies are at different development phases, they can solve particular problems in the processes and operation.
  • Building capabilities proactively, both technical and soft, and managing talent. Companies can tailor the skills training programs to the entire organization, from front-line worker to managing director, based on their digital affinities. Meanwhile, building an agile architecture for organization and performance management can support intelligent transformation and technology innovation as well.
Actually, manufacturing in some countries, such as the United States, Germany, Japan and South Korea, is also facing similar purgatories as the manufacturing in China. Therefore, the recipe of Foxconn Industrial Internet has the potential to be scaled to global manufacturing companies.
III IIoT intelligence to support manufacturing transformation
IIoT intelligence can support the construction of three pillars and two capabilities required by manufacturing transformation, which are explain in Section II-B. In the following subsections, we will explain the connotation of IIoT intelligence and its effects on smart manufacturing.
III-A Understanding IIoT intelligence
We give our definition for IIoT intelligence:
IIoT intelligence refers to a series of techniques, methods, productions and platforms that can been taken throughout the entire value chain to build capabilities of digital connection and perception, intelligent analysis and cognition, real-time decision-making, etc.
The value chain of manufacturing involves research and development (R&D), production, operation and maintenance, marketing, management, services, etc. All segments of the value chain are the launching scenarios of the IIoT intelligence. That is, the essence of IIoT intelligence is to realize innovative industrial applications, such as digital R&D, efficient and immediate decisions, rapid line reconfiguration, through in-depth combining intelligent technologies with industrial scenarios, mechanisms and knowledge.
IIoT intelligence is improving by iterating continuously. At the beginning, data throughout the value chain is collected and some machine learning algorithms are used to learn specific knowledge to all segments of the value chain. Then, based on the data and knowledge, optimized models and algorithms are designed and verified. Afterwards, the optimized solution is scaled within the value chain. Finally, new data will be generated, a new round that improves the value chain will start.
III-B How does IIoT intelligence achieve impacts in manufacturing ?
IIoT intelligence can provide some innovative solutions for segments of the value chain of manufacturing, and support to build the pillars and capabilities (shown in Fig. 3). The solutions are able to improve direct labour productivity and employee engagement, raise product quality and overall equipment effectiveness (OEE), reduce R&D costs and energy consumption, etc. We present a close look at two cases to demonstrate the idea as follows.
Figure 3: The secrets to intelligent transformation of manufacturing
III-B1 Digital experiments to reduce product R&D costs
Traditionally, the R&D of industrial products should experience five steps: project approval, product research, production preparation, process optimization, and declaration of announcement catalogue. At the stage of product approval, the product development mission statement and R&D budget are required, and the documents need the approval of the president. While researching and developing new products, design drawings and documents of products are required, and trial production of products with some success is expected as well. The production preparation aiming at putting forward production technologies and process planning, and piloting the solutions, according to the design drawings and documents of products. Missions of the process optimization are improving initial process planning, organizing production, and discussing employee income. In the phase of declaration of announcement catalogue, the product declaration planning should be determined according to the product development specification and sales plan.
Revising the processes of the R&D of products, we find several rounds of trial production, testing and optimization in the physical production lines are important for achieving a new product. However, each round of the physical experiment costs a lot of manpower, material, money and time. Therefore, the traditional R&D process is time-consuming, expensive and environmentally unfriendly.
IIoT intelligence can reduce costs of product R&D by transferring part of the physical experiments into digital world. Specifically, IIoT intelligence is able to take full advantage of knowledge from historical experience and data, then map industrial processes into the digital world, where some trial and error experiments can be carried out at low costs. Besides that, some faults are possible to be predicted automatically according to industrial knowledge, mechanisms and historical experience in the digital world. Some improvement suggestions are likely to be given by some recommendation technologies. Receiving the detected/predicted faults and some valuable ideas, the project team is able to immediately revise the designed products and optimize the architecture of new products.
III-B2 Remote monitoring and control to increase labour efficiency
Traditionally, workers are expected to physically participate in production or monitor equipment operation in the workshop, in order to detect and correct errors during production in time. The maintenance of production equipment depends on the on-site service of technicians.  The deployment of production lines also relies on manual analysis and expert experience.
IIoT intelligent connection and perception technologies can connect equipment, workers and intermediate products in the workshops. It enables remote monitoring of production conditions and remote control of manufacturing equipment, which improve safety through unmanned operation and reduce the number of workers on dangerous sites. The technology can also help improve OEE (overall equipment effectiveness) and labour efficiency by boosting one-man-one-control to one-man-multi-control. With the help of remote monitoring and control, managers are able to flexibly schedule workers according to the progress on-site. In addition, the technologies introduce some white-collar jobs, allowing workers can work in a comfortable office. In this way, employment willingness to the manufacturing industry can be improved and the problem of labour shortage in the manufacturing industry can be alleviated.
III-C The overall effect of IIoT intelligence on smart manufacturing
In summary, IIoT intelligence can promote the rapid development of smart manufacturing from the perspectives of functions, operation, deployment and application scenarios.
  • Functions: IIoT intelligence can help companies connect human-cyber-physical spaces, collect industrial data synchronously and accurately, build effective data analysis models, propose scenario-oriented optimization methods, develop microservice architecture, etc.
  • Operation: IIoT intelligence builds cloud manufacturing business model for enterprises to enhance returns [33, 34, 35]. Cloud manufacturing create intelligent factory networks that encourage effective collaboration, through encouraging enterprises to encapsulate distributed manufacturing resources into cloud services and managing them in a centralized way. Clients can use and pay for the cloud manufacturing resources according to their requirements.
  • Deployment: IIoT intelligence provides the cloud-edge-terminal architecture to deploy computing resources in a cost-friendly way. Except that, IIoT intelligence can give efficient production lines deployment suggestions to improve factory output.
  • Application scenarios: IIoT intelligence can be used to solve specific problems in some application scenarios, such as predictive and remote maintenance, product and equipment monitoring, product design feedback optimization, product process traceability, renting manufacturing capacity, product quality inspection, etc.
IV Hierarchical development architecture of IIoT intelligence
To build a systematic understanding of how IIoT intelligence deals with issues in the transformation process of smart manufacturing, this paper presents a hierarchical development architecture for IIoT intelligence. The architecture consists of equipment, network, software, model and optimization from bottom to top, as shown in Fig. 4. The classic architecture of IIoT usually include the bottom two layers [8, 9, 11], however, it cannot bear the weight of the various intelligence that have emerged in recent years. Therefore, we extends software layer, modeling layer, and analysis and optimization layer over the classic architecture.
The functions of each layer in the hierarchical architecture are as follows:
  • Equipment Layer: Technological innovation in the equipment layer builds the basis of automatic production. Especially, industrial robots paired with smart sensors will replace human labour to achieve automatic production to some extent, or assist workers engaging in flexible production. Cloud/Edge/fog computing ensures sufficient computing resources to support the development of IIoT intelligence.
  • Networking Layer: Communication technologies in the networking layer constructs a network that supports high reliability and low delay transmissions of human-cyber-physical information flows. The network can overcome isolated islands of information in the companies, which will greatly accelerates the development of smart manufacturing.
  • Software Layer: Industrial software provides digital expression of industrial processes by collecting and mining industrial data. Industrial software can meet the requirements of enterprises to strengthen control over various industrial processes, making industrial processes transparent, and also conducive to optimizing production and management.
  • Modeling Layer: Modeling technologies in the modeling layer are expected to accurately model industrial processes and scenarios, based on the digital expression from industrial software. At the same time, the model layer should support the analysis and optimization layer via providing the necessary data, model and management control interfaces.
  • Analysis and Optimization Layer: The analysis and optimization layers consists of some specific algorithms. The algorithms are designed to solve specific problems in the industrial processes based on the industrial big data coming up from the bottom and models built in the modeling layers.
Figure 4: The hierarchical development architecture of the IIoT intelligence. The equipment layer builds the basis of automatic production. The network layer connects human-machine-things, and overcomes isolated islands of information. The software layer provides digital representation of industrial processes. The modeling layer makes digital twin for physical processes to connect virtual and physical spaces. The analysis and optimization layer is responsible for mining industrial big data and optimizing industrial processes.
The transmissions of information flow in the hierarchical development architecture are bidirectional.
  • The data is transmitted from bottom to top. The industrial equipment and workers produce data constantly. The networks ubiquitously connecting human-machine-things transmit data to computing devices with specific industrial software. After collecting and roughly mining data, software provides required information to modeling layer, so that accurate models can be constructed. Finally, the information and models are used by the specific algorithms in the analysis and optimization layer.
  • The data is transmitted from top to bottom. An optimized solution is firstly issued in the analysis and optimization layer, and then assigned to modeling layer. After modules in the modeling layer understand the solution, they will inform corresponding management and control software. Finally, the equipment layer will carry out the solutions.
In the following subsections, the motivation and connotation of each layer in the hierarchical development architecture are introduced in detail.
IV-A Equipment layer
IV-A1 Motivations
Facing the crisis of labour shortage and high labour cost, the manufacturing industry’s desire for a machine that can work like or even better than humans has become increasingly strong, thus promoting the invention and innovation of industrial robots. Except that, there are other reasons that motivate enterprises [36]. a) There are many dull, dirty, dangerous and/or delicate tasks that can be done by robots, which can improve health, safety and job satisfaction of employees. b) The rapid development of robotics technology is continuously improving the automated and flexible production capacity of robots, while also reducing the procurement and maintenance costs. c) Modern robots will support a smaller carbon footprint [37][38], which is profitable for building low-carbon factories according to national requirements. d) More and more robots support plug-and-play integration and their control programming becomes much easy, which allow operators to become proficient in using robots with general training [39]. e) Human-robot collaboration will strengthen flexible production [40][41]. In addition, the aging society of the world has also released the demand for liberation from manual labor. Under the circumstances, governments also propose some important policies111details of the policies are summarized in https://ifr.org/ifr-press-releases/news/investment-in-robotics-research-global-report-2021 to promote development of industrial robots, such as Made in China 2025, the New Robot Strategy In Japan, The Intelligent Robot Development and Supply Promotion Act of Korea, The new European Framework Program Horizon Europe,Germany’s High-Tech Strategy 2025, The National Robotics Initiative (NRI) in the USA.
Except robots, in order to achieve precise perception of the state of the production environment and high-precision control of the production process, it is necessary to introduce smart sensors to improve five senses222taste, touch, smell, sight, hearing of industry. In addition, with the deepening of industrial digital transformation, a large number of industrial data storage and analysis tasks have emerged. To complete these tasks, massive computing and storage resources are required.
IV-A2 Connotation
Equipment in the equipment layer can be divided into two categories, that is, production equipment, which includes industrial robots and sensors, and computation devices, which include computation resources and storage resources on cloud/edge/fog.
Industrial sensors perceive a physical quantity at the input and send a corresponding electrical signal at the output. There are a lot of industrial sensors that can be used to improve five senses of industry, including air quality sensors, ambient light sensors, distance sensors, image sensors, industrial motion & position sensors, infrared detectors, industrial pressure sensors, speed sensors, etc. Industrial sensors can be found in nearly every modern manufacturing process, regardless of complexity. As a result, it is nearly impossible to conceive an automated system without sensors [42].
An industrial robot refers to a robot system that can be used for manufacturing. Industrial robots are automated, programmable and capable of movement on three or more axes. Industrial robots are gradually replacing human to overtake dangerous and delicate tasks. Typical application that robots can realize include welding, painting, assembly, disassembly pick and place for printed circuit boards, packaging and labeling, palletizing, product inspection, and testing. Moreover, with the help of smart sensors, robots can ensure accomplish the tasks with high endurance, speed, and precision.
Cloud computing that brings computing services, including servers, storage, databases, softwar, etc. over the Internet, has become a economic and flexible computation resources deployment solution for enterprises, in recent years. Most cloud computing services fall into four broad categories: infrastructure as a service (IaaS), platform as a service (PaaS), serverless, and software as a service (SaaS) [43]. However, because cloud computing that resources are usually far away from the data sources, brings a long latency for IIoT data analysis, edge computing and fog computing that bring computation resources near data sources offer new ways of processing and analyzing data in real time. Currently, cloud-edge-fog architecture has become an effective architecture to provide sufficient computation and storage resources that are required by IIoT intelligence.
Figure 5: The networking helps to realize smart factories from information islands to collaborative operations
IV-B Networking layer
IV-B1 Motivations
The revolution of IIoT networks is inspired by global trends[44]. Firstly, because of more and more smart equipment and staff accessing the network in the domain of smart manufacturing, and the connection between systems, human, processes, locations and devices getting more distributed and complex, the economic value of the network for enterprises are growing. Secondly, the inherent unpredictability of business that the network quickly corrects the transmission modes to adjust to the new services, processes and models. Thirdly, many manufacturing processes are time sensitive and mission critical, which requires the communication time of network to be more precise and predictable[45]. Moreover, advanced networks are expected to offer the promise of the environmental sustainability.
Technology improvements encourage IIoT network revolution as well. The applications and data of enterprises are leaving the premises and moving to the public/private clouds. Part of the monolithic applications in many cases are being modularized into microservices that are delivered via a variety of virtual and physical workloads across the entire enterprises. Besides that, the APPs are built and updated continuously and rapidly to meet the demands of business modes. Under the background of applications and microservices revolution, the network is expected to be a growing set of nerve clusters that can be anywhere along the cloud-edge-terminals. The network should understand how innovative application models work and dynamically extend the applications to wherever the applications are hosted.
CISILION summarizes four primary objectives for the new network in terms of aligning to the business, abstracting complexity, assuring performance and reducing risks [44]. The network in IIoT intelligence is expected to achieve some similar characteristics. As presented in Fig. 5, which refers from [46] and Prediction from TF security [47], the network connecting human-machine-things helps overcome the serious problem of information island, provides high-speed, wide-converage and low-latency communication, and supports the development of overlay applications, such as industrial big-data mining and immediate decisions. To be specific, in the increasingly demanding manufacturing environment, the network needs to be able to adapt quickly to changing manufacturing requirements. The network should provide convenient access of fast-changing set of equipments, staff, applications and services, so that resource scheduling and management can be focus more on. Providing steady continuous service without interruption is also essential for the network used in the manufacturing.
IV-B2 Connotation
All feasible networking methods that connects human, sensors, robots, cloud/edge/fog computing resources and other smart devices belong to the networking layer of IIoT intelligence.
Some advanced networks are proposed in recent years. Representative progresses involve software-defined network (SDN), white box switch, information centric network (ICN), deterministic networks, time sensitive networks, 5G and network slicing, etc. These achievements are driving the network to realize high performance, including rich bandwidth, high-speed and steady transmission, massive connectivity, and so on. Here some techniques are briefly introduced as examples to illustrate the effects of advanced technologies for the future IIoT network. SDN exploits the ability to decouple the data plane and the control plane in routers and switches [48], which enables the network to be intelligently and centrally controlled and managed, regardless of the complexity of the underlying network technologies [49]. ICN is a promising candidate for the paradigm of the future internet [50]. Inspired by the fact that it is the information dissemination rather than pair-wise communication between end hosts, ICN switches the IP-based Internet to a information based architecture by in-network caching and storage [50][51]. The fifth generation (5G) wireless communication aims to achieve revolutionary leap forward in terms of latency, massive connectivity, network reliability and energy efficiency [52][53]. The network slicing is an unique technique that is firstly used in 5G to achieve the ambition of 5G. Network slicing facilitates multiple logical self-contained networks on top of a common physical infrastructure platform, so that 5G is able to support very diverse and sometimes extreme requirements in an isolated and transparent approach [54]. Some industrial instances that network slicing is used are given in the work [55].
IV-C Industrial software layer
IV-C1 Motivations
In an increasingly digital manufacturing, human, smart sensors and robots are generating industrial data continuously, and networks are busy for transferring data. The foundation of IIoT intelligence has been built by equipment layer and networking layer. However, there is a lack of software that can be used to manage data and production process.
Figure 6: The primary functions and application instances of R&D software, production control software, information management software and embedded software. The grey blocks represent the expected effectiveness of the industrial softwares. The yellow blocks show the corresponding functions. The orange blocks include representative instances.
IV-C2 Connotation
The industrial software is the product of the integration of industrial knowledge and information technologies. The product is updating iteratively as the long-term accumulation and application of industrial innovative knowledge. According to the functions and characteristics, the industrial software can be divided into four categories: R&D [^3] Research and Development software, production control software, information management software and embedded software[56]. Fig. 6 presents the primary functions and application instances of the four types of industrial softwares.
The industrial softwares help managers enhance control for the whole industrial processes and optimize production and operation schedules. For example, INTMES [^4] http://www.intmes.com/, whose primary business is designing personalized software solution for companies, provides a customized MES/SCM software555details shown in http://www.intmes.com/solution-content-CASE-2.html to Hanteng automobile Co.Ltd, a vehicle manufacturing company. With the help of MES/SCM, Hanteng’s workshop supports transparent management and fine control, and its labour efficiency is improved as well.
The industrial software is highly complex for some reasons. First of all, there are various industries, among which process industry and discrete industry contain 39 major categories, 191 medium-granularity categories and 525 fine-granularity categories. Secondly, many professional knowledge and experience in machinery, electronics, optics, acoustics, sound control, fluid heat treatment are required to support research and development of industrial softwares. Thirdly, the complexity of industrial products is wide-span, from simple products such as clothing and toys to complex products in aviation and satellite. Moreover, the industrial softwares engage in all stages of industrial processes, while the industrial process is complex, including R&D, production, marketing, operation, maintenance, supply chain management and others. Deployed equipments are also various, functions involve production, implementation, test and measurement. Then, it is difficult for manufacturers to cooperate. For example, Boeing747 is jointly produced by more than 16000 enterprises with different scales from six countries. Finally, industrial software has high requirements for real-time reliability, because unreliable data is possible to cause huge problems in terms of cost and safety.
In recent years, governments are promoting the rapid development of industrial software by policy incentives, at the same time, innovative technologies also provides opportunities to update industrial software. As the industry environment becomes more and more competitive and complex, the industrial software are upgrading in three ways, including systematic platform, cloud-based deployment and lightweight development. To be specific, early-stage industrial software are oriented specific scenarios, such as CAD for designing, CAE for simulation, ERP for enterprise management, etc. Current industry requires collaboration among multiple companies. Therefore, some systematic platforms are expected to support the cross-company interaction. The systematic platforms should integrate and utilize resources and tools to provide solutions for smart manufacturing. On the other hand, developing cloud-based industrial softwares [57][58] is an effective approach to deal with current operation models (IaaS, SaaS, PaaS) and customization demands. Guaranteeing data interconnection, real-time interaction and remote monitoring [21] is important topic in smart manufacturing, which promotes to develop lightweight mobile industrial APP.
IV-D Modeling layer
IV-D1 Motivation
In the architecture of IIoT intelligence, the modeling layer should achieve two functions, that is, modeling physical processes into the digital spaces, at the same time, transferring decisions made in the digital space to physical world. The development of modeling methods are driven by future intelligent applications, such as remote control, predictive maintenance, process optimization and so on. Specifically, in order to perform well on these scenarios, it is necessary to model particular objects accurately, and it is better to maintain uninterrupted connections with the objects to update the models in real time and transfer instructions.
IV-D2 Connotation
Modeling methods that can build connections between physical world and cyber world f or IIoT belong to the modeling layer.
The digital twin is the most promising technology to realize the interactive between cyber spaces and physical spaces. The representation model of digital twin is shown in Fig. 7. The digital twin has five necessary keypoints: data, connection, model, interaction, and application. Consequently, the digital twin can run through the entire life cycle of physical processes, and provide capabilities with description, analysis, diagnosis, prediction, prescription, cognition, and so on [59][60]. Currently, the industrial-research-university communities explore the digital twin in two aspects, modeling and simulation. Modeling refers to establishing 3D virtual twins for physical entities. The twins are expected to reproduce appearance, geometric association, motion dynamics, and other attributes of the physical entities. The simulation is an activity to calculate, analyze and predict the future state of physical entities based on the virtual twins with industrial knowledge, mechanism, and data. As far as researches contributions at this stage, the modeling and simulation of digital twin depend on industrial software, including CAD, CAE, etc.
Figure 7: The representation model of digital twin
Besides the digital twin, there are some other methodologies that can model certain industrial scenarios, such as Use Case Maps (UCMs), message sequence chart (MSCs), digital shallow, etc. The UCMs is a real-time system oriented and scenario-based modeling method [61][62]. The UCMs uses abstract symbols to present the structure and interactive behaviors of complex systems, so that developers can understand the overall system easily. The UCMs provides an intuitive and clear visual system description method, which can realize flexible conversion between formal and informal description of the system. The MSCs, also known as timing diagrams, message flow diagrams, are effective graphical and textual languages used to describe the system messaging and are widely used in software development practices. There are three popular MSC software versions, namely MSC-92, MSC-96 and MSC-2000 respectively. The MSC-92 standard refines the basic concepts in the system and uses formal expressions to describe the basic processes. In the MSC-96 version, structured language is added compared with the previous version. The MSC-2000 version has made improvements in four aspects: control flow, quantitative representation time, data description and conditions.
IV-E Analysis and optimization layer
IV-E1 Motivations
With the help of modeling layer, the connection between cyber space and physical space is constructed. However, the connection only provide a channel to control physical processes based on the optimization in the cyber space. It cannot solve specific problems in the industrial environment, such as manufacturing processes planning and schedule, topology optimization, and prognosis, etc. Therefore, we design the analysis and optimization layer to support the development of IIoT intelligence in specific problems and scenarios.
IV-E2 Connotation
intelligent algorithms that can solve specific problems in the industrial environment belong to the analysis and optimization layer. There are too many scenarios and problems to enumerate them all. Due to space limitation, this paper only identifies some issues of greater concerns, including manufacturing processes planning and schedule, topology optimization, and prognosis.
Manufacturing process planning and scheduling
Manufacturing process planning determines how a product will be manufactured by selecting and sequencing manufacturing operations. The process planning aims at improving some specific performance (e.g., shorten processing time, minimize production cost) while satisfying a set of domain constraints [63] [64]. The manufacturing scheduling is the process of assigning manufacturing resources over time to the set of manufacturing processes in the process plan [63]. It is an optimization problem that schedules of resource allocation will be proposed with respect to criteria such as delay, throughput, or cost. The process planning and scheduling problem is becoming increasingly important to improve the flexible manufacturing. However, the problem is typically NP-hard, that is, it is impossible to get optimal solutions without the use of an essentially enumerative algorithm, while the computing complexity grows significantly as the problem scale increases.
Computer-aided process planning (CAPP) [65][66] based approaches are useful for the problems, such as petri-net-based approaches and knowledge-based expert system, agent-based methods, and heuristic algorithms. Petri-net tools can also be used to analyze the performance of the system, generate code, simulate the system and perform model checking on it [67, 68, 69, 70]. Knowledge-based system provides necessary expert information, featuring information input, recovering comparable plan and making essential alteration for manufacturing process planning and scheduling [71, 72, 73]. Agent technology [74] has received significant attention and achieved important progress, both in terms of algorithms [75, 76, 77] and platforms [78, 79, 80, 81]. Cooperative intelligent agents are also used in developing distributed CAPP systems for process planning and scheduling problems in manufacturing [82, 83, 84, 85]. Heuristic approaches [86][87], like climbing, simulated annealing, Tabu search, and GAs, try to replace the exhaustive search strategies with some sophisticated experience.
Topology optimization
Topology optimization is a mathematical method that answers the question of how to place material within a prescribed design domain in order to obtain the best structural performance [88]. The topology optimization happens at the initial stage of the product manufacturing process [89]. It is efficient because automated optimization processes are employed to generate the conceptual designs instead of the conventional trial-and-error approach [90]. Some IIoT intelligence technologies are used to maximize the performance. Surveys of representative manufacturing-oriented topology optimization approaches [90][88] summarize the efforts to parameterize the topology (SIMP method [91][92], level set method [93], spline based methods and some limitations), machining-oriented topology optimization methods (length scale control [94] and geometric feature based design [95]), and injection casting oriented topology optimization (thickness control [96]).
Industrial prognostics
The researches about industrial prognostics mainly focus on estimating and anticipating events of interest regarding industrial assets and production processes [97]. Particularly, the industrial prognostics are regraded as a data-based workflow in the work [97], and two important research topics are mentioned. $(a)$ The descriptive prognostics that mainly summarize data and infer insights from the data. Consequently, the gained information from data can help detecting events of interest or to estimate the health status of the industrial asset, product or process under study. Important research topics are pattern recognition and classification [98][99], and health management [100][101]. $(b)$ The predictive prognostics utilize a variety of data fusion, statistics, modeling, and machine learning techniques to study recent and historical data, to learn prognostic models, which make accurate predictions about the future status of the monitored asset. The condition-based maintenance [102][103] and the predictive maintenance [104][105] are two of the interesting topics.
Some prognostics approaches are summarized in the survey [106], such as $(a)$ physics-based approaches [107] achieve prognostics by using mathematical models to describe the degradation mechanics or damage propagation of a machinery component. $(b)$ Data-driven approaches [108][109] use the collected data (usually condition monitoring data) to map the characteristics of damage/degradation state to achieve prognostics. $(c)$ Hybrid approaches [110] combine the advantages of different approaches to improve the prediction accuracy or extend prognostics model applications.
TABLE I: Digitization transformation of 4 lighthouse factories with IIoT intelligence
AGCO/Fendt, Marktoberdorf, Germany
Henkel Laundry and Home Care, Germany
Johnson & Johnson Vision Care, Jacksonville, Florida
Weichai Power Company, Weifang, China
Analysis and Optimization Layer
  • Advanced big-data analysis and quality monitoring
  • Integrating data evaluation and optimization tools
  • Standardization of measurements and calculation of KPIs
  • Adaptive process control
  • Advanced analysis algorithms to connect manufacturer, patient, professional opticians and retailer
  • Big-data analysis, remote maintenance
Modeling Layer
  • Digital twin
  • Digital shadows of 33 sites and 10 distribution centres
  • Augmented-reality tools for testing product comfort
  • Digital rapid modelling design
  • Virtual development simulation
  • IIoT-based testing
Software Layer
  • Holistic manufacturing execution system (MES)
  • Laundry data foundation
  • Henkel laundry digital backbone
  • An adaptive flexible modular platform (FMP) for rapid new product introduction
  • An interconnected visual-value-chain control tower
  • A mobile/web platform
  • An end-to-end product development system
  • An IoV system to collect real-time data on engine speed, fuel consumption and power under various working conditions
  • An app-based car networking platform provides users with efficient service solutions
Networking Layer
  • Iot realizing a consistent data platform across all systems
  • A digital backbone connecting factories to a central data repository in order to collect real-time energy consumption data
  • IIoT-enabled and fully validated closed loop
  • Intelligent IIoT-based testing
Equipment Layer
  • Manufacturing-related robots and sensors
  • 3500 interconnected sensors are used to measure electricity, fossil fuel, water, compressed air, steam consumption
  • Advanced robotics and intelligent material handling
  • Manufacturing robots and increased sensors are used to ensure real-time acquisition and uploading of test results
Returns
  • 60% lower cycle time
  • 37% assembly line volume increase
  • 30% reduction in time-to-identify quality issues
  • 36% lower sustain ability footprint
  • 10% reduction in processing costs
  • 25% reduction in logistic costs
  • Reducing Henkel’s energy consumption by 38% and its water consumption by 28%
  • 30% faster time to market on new products
  • Maximizing yield, reducing downtimes
  • 11% OEE improvement
  • Enabling a 50% reduction in the time to scale worldwide volume of new products
  • New product development is shortened from 24 months to 18 months
  • The product design reuse rate has increased by 30%
  • Reducing test costs by more than 20%
  • Reducing labour costs by 75%, shortening the R&D cycle by more than 20%
V A close look on lighthouse factories with IIoT intelligence
In this section, a close look on lighthouses and collaborations across companies is presented to demonstrate that IIoT intelligence is able to empower intelligence transformation and upgrading to smart manufacturing, and augment the confidence to explore and exploit IIoT intelligence.
Since 2016, the World Economic Forum, in collaboration with McKinsey & Company, has been monitoring progress in advanced manufacturing worldwide. In 2017, the World Economic Forum identified 40 initial advanced manufacturing use cases that began to pilot intelligent technologies with some success. In 2018, 16 companies were recognized as leaders that achieve step-change results, both operational and financial, across individual sites [28]. In 2019, 28 additional facilities were identified as end-to-end lighthouses [29].
These advanced companies succeed in transforming factories, value chains and business models to generate compelling financial, operational and environment returns [29]. The specific key performance indicator (KPI) improvements are reflected in factory output increase, productivity increase, overall equipment effectiveness (OEE) increase, quality cost reduction, waste reduction, energy efficiency, change-over shortening, etc. These impressive progresses depend on expanding IIoT intelligence within the manufacturing plant environment, for example, automation, digitization, big-data analysis, remote control, 3D printing, etc.
According to the inside perspectives from the global lighthouse network, this paper closely looks the contributions of IIoT intelligence for smart manufacturing. This paper selects some lighthouse companies, and summarize the improvements from the perspective of hierarchical development of IIoT intelligence. The upgrade behaviours and corresponding returns are presented in Table I.
Besides the successful end-to-end digital transformation lighthouse factories, some other collaborations are gradually released. Some important news are declared. Reports of ”How 5G Will Revolutionize Manufacturing” and ”Powering industries and a smart world with 5G, Cloud & AI” are published[111][112]. In 2019, Haier, China Mobile, Huawei and Mstar released the world’s first 5G smart factory joint solution [6]. The factory implements intelligence in terms of quality detection, maintenance, machine collaboration, material transportation, enery management and campus security under help of IIoT intelligence. In 2021, Midea Group, China Unicom, and Huawei rolled out Guangdong’s first 5G fully-connected smart manufacturing demonstration site at the factory of Midea’s Kitchen Appliance Division in Shunde district, Foshan city [7]. In the project, IIoT intelligence will enable achievements in multiple aspects. Specifically, the high-performance 5G dedicated network supports highly efficient production on the production line; the AI brain monitors and predicts robots’ behaviors and ensures that they are unified; machine vision is adopted to facilitate product quality inspection, significantly reducing the need for manual inspection; and the AGV carts deliver materials to specified locations.
VI Technology studies on IIoT intelligence
There is no doubt that the rapid and effective development of IIoT intelligence cannot leave some crucial technologies. Because of the space limit, we identify some of the most influential technologies for IIoT intelligence, and clarify their contributions in the development of IIoT intelligence. The application progresses of these technologies are described as follows.
VI-A Industrial robots as intelligent labors
The industrial robot usually refers to the manufacturing-oriented multi-joint mechanical arms, or other mechanical devices with multiple axes. The primary characteristics involves automatically control, programmable, multi-purpose, fixed in place or mobile, etc. Based on the mechanical structure, the robots can be classified into six types, including articulated robot, SCARA robot, cartesian robot, Delta robot, cylindrical robot and others [113][23]. This paper presents some widely used robots. The articulated robot refers to a robot whose arm has at least three rotary joints. It has a large operating range, supports flexible movements and can grab objects close to the fuselage. The cartesian robot is a robot whose arm has three prismatic joints and whose axes are correlated with a cartesian coordinate system. The cartesian robots have simple kinematics, high positioning accuracy, and great stability. But its flexibility is poor, and the operating range is small. The Delta robot is a robot whose arms have concurrent prismatic or rotary joints. Its motor is mounted on a fixed base. By this way, the inertia of the robot in the process of motion can be reduced.
The industrial robots are mainly used to replace human in automatic and/or flexible production to some extent [22][23]. According to the statistics of operational sotck of industrial robots by customer industry and application - world made by IFR(the International Federation of Robotics)[36], the largest application of industries of industrial robots is automotive manufacturing, following that electronics is important as well. As a consequence, the automatic production and flexible manufacturing of these two areas are developing rapidly [31][28]. In terms of usage, most industrial robots are used for handling, followed by welding and assembling. These tasks usually require lots of physical energy and part of content has bad impacts on health.  At the same time, the statistics about annual installations also demonstrate that industrial robots have played a significant positive role in promoting industrial processes, and there is still large room for further development.
The most three important components of industrial robots are reducer, servo motor and controller. The reducer is the kernel of industrial robots, the servo provides energy for robots, and the controller likes the brain of robots. The tree primary components occupy 35%, 20% and 15% respectively, accounting for 70% in total. To further improve performance of robots, researchers and developers should break through technical barriers in these three parts.
For industrial robots at current stage, a relatively complete industrial chain has been constructed and each stage has leading enterprises. Except the basic materials, like metal, non-metal and basic components, the techniques of making three primary components are handled in some companies, such as, Siemens, Yaskawa, Eston, Nabotsk, Hammer, KEBA, ABB, etc. After getting enough materials and primary components, some companies focus on integrating robots. Related leading enterprises include Fanuc, Yaskawa, Nizhi, Eston, and so on. In addition, some well-known companies, such as Siemens, Fanuc, ABB, etc., also overtake business of system integration, which refers to deploy production lines using the robots and some software.
VI-B Machine vision system as eyes of industries
TABLE II: Comparison between human eyes and machine vision
Human Eyes
Machine Vision
Adaptability
Strong
Relatively strong, less demanding on the environment
Speed
The visual persistence of 0.1 second makes it impossible for human eyes to see faster moving targets
The shutter time can reach about 10 microseconds
Precision
Low, unable to quantify
High, micron accuracy, easily quantified
Grayscale resolution
Low, only 64 gray level resolution
High, currently 256 gray levels are generally used, and the acquisition system can getgray levels such as 10bit, 12bit, 16bit, etc.
Repeatability
Low, easily affected by emotions, physical conditions, etc.
Always uniform methods and standards
Data collection and the possibility of future intelligence applications
No, the human eye does not have the ability to analyze after collecting data
Strong, with strong data collection and analysis functions, it is a huge data entry at the edge layer
The goal of a machine vision system [114][115] is to create simplified models of the real world based on sensed images. These models should be understandable to computers and can restore real information (geometric shape, color, texture, category, printed matter, etc.) from images. Compared with human eyes, machine vision has significant differences and benefits, as presented in Table II. Thanks to these advantages, machine vision systems have been applied in industry.
Machine vision systems have great differences between various industrial tasks. For general practical applications, there are some shared common performance requirements in terms of precision, speed, environmental robustness, and easy to use [116]. Without accurate measurement, the machine vision cannot successfully guide the robotic hand to perform automatic assembly of smartphone or mechanical watch  [117, 118, 119]. In the high-speed assembly, the average time of template matching is less than several microseconds for electronics manufacturing [120, 121]. A brief overview about how machine vision augments manufacturing technologies is given as follows.
The manufacturing techniques, including machining (milling, grinding, polishing), welding, injection molding, etc., can be improved with an auxiliary machine vision system. As an efficient and intelligent machining of complex free-form surface, robotic grinding demonstrates flexibility, intelligence and cost efficiency under the guidance of machine vision [122][123]. The raw workpiece is located by matching the measured point clouds and corresponding theoretical model. Thus, the machining error and allowance distribution for free-form surfaces can be optimized. Due to the accuracy requirement of grinding, accurate point clouds scanning and accurate calibrating between the machine vision system (scanning system) and the robot system are key to achieving high performance robotic grinding.
Equipped with machine vision system, the intelligence robotic welding is drawing attention from industrial community due to the high quality of product [124]. Robotic seam tracking is challenging due to the strong noise images(reflection, strong arc light and the large spatter), and a convolution neural network is proposed to extract the feature points of weld seam image [125]. A machine vision approach is introduced to tracking the welding electrode tip for assessing welding arc stability and polarity in manual metal arc [126]. The machine vision system thus becomes essential for welding monitor, diagnosis, prediction, and control. How to obtain useful information from noisy sequential images and how to use such information to improve the welding skill are two open questions for intelligent robotic welding [124].
With the help of machine vision, the robotic assembly has a long history due to the efficiency [127][128]. With the increasing capability of machine vision, the smart assembly demonstrates flexibility as well as challenges [129]. One challenge of robotic assembly is that the positional error of machine vision must meet the minimal requirement of individual task. Otherwise, the automatic assemblies may fail [130][131]. One direct solution is to improve the positional accuracy of machine vision by micro-measurement, such as eye-in-hand visual system [118] or optical configure of micro-view [132]. To achieve successful assembly, a data fusion trend with multi-modal sensors is considered by several recent works [133][134]. Regarding the automatic assembly, the challenges of machine vision are the accuracy requirement, multi-modal sensor fusion, deformable object manipulation, and advanced human-robot interaction.
Defect detection is highly recommended to ensure product quality following every manufacture/assembly procedure. The machine vision system is popular in defect detection due to its efficiency and attractive cost. The geometrical surface defect inspection is often achieved by point clouds registration between the scanned points and its template [135][136]. While, the assessment for compliant parts is still an open challenge [137][138]. The assembly defect detection, called as assembly line quality control, is to check the assembly correctness [139][140].
The on-line visual feedback of machine vision measurement (position, shape, velocity, etc.) enables the machining, welding, injection molding techniques with improved manufacturing quality and speed. Regarding the manufacturing of mechanical parts, the challenges of machine vision are the high accuracy requirement, harsh imaging conditions (occlusion, unwanted light, cutting chips, and cutting fluids), and field measurement of deformable/fluid materials.
VI-C Networking facilitates ubiquitous connectivity
Current network is the product of the deep integration of the new generation of information, communication and modern industrial technologies. It has been one of the magic weapons to realize digital transformation of manufacturing. This paper will briefly introduces some related progresses of typical network technologies and applications, such as 5G, time sensitive networking, and identity resolution.
The 5G network can connect autonomous industrial machines, ground vehicles, robots and software in an URLLC wireless communication, aiming at assistance of real-time control and management of industrial machinery towards smart factories [52][53][141]. The authors in [141] study 5G-enabled industrial Internet of Things, considering using machine learning to handle complex tasks at industrial machinery and 5G networks management,configuration,and control. Moreover, the authors design transfer learning based systems for Industrial Internet of Things (IIoT), and discuss how transfer learning can overcome the demand for high-quality large data samples required to training neural network in IIoT. An intelligent and green resource allocation mechanism for the IIoT under 5G heterogeneous networks is proposed [142], so as to ensure the high reliability and energy efficiency of IIoT services. Networking slicing is an essential technique to support massive connectivity and service isolation for 5G, so as to realize scale-up applications in the industry [54][55].
Time-Sensitive Networking (TSN) [143] is a standardized technology that provides deterministic information transmission on traditional Ethernet. In [144], the authors introduce Open Platform Communication Unified Architecture TSN (OPC UA TSN) as a new technology and present the current view. In [145], the authors present an approach to integrating heterogeneous Industrial Ethernet applications over TSN, and deploy different real-time schedulers that are embedded in a Software-defined Network (SDN)[48][49] controller to allow different applications with various traffic patterns and Quality-of-Service (QoS) requirements to share the same network. In [146], the authors introduce the latest research progress of industrial communication systems, and analyze the technical challenges and future trends for TSN.
The identity resolution as an important network technology [147] builds the nerve endings of networking while used in industry. The kernel functions of the identity resolution are distribution, registration, management, analysis and routing of the identity of goods. Meanwhile, identity resolution can support the whole life cycle management of equipment, people and materials. It not only helps realize data inter-operability and data mining, but also provides a necessary condition for the intelligent management of manufacturing. There are many contributions about identity resolution systems [148, 149, 150]. Among which, RFID (radio frequency identification) based methods have made great progresses. RFID is an automatic object identification and data collection technology by using radio waves. Detail development of RFID from tags to middleware and application is discussed [151]. An insight to the application of RFID technology in the management of manufacturing and supply chain is summarized [152]. Some challenges in implementing the technology and potential future manufacturing application are discussed in the work as well [152]. With consideration of the information accuracy, time consuming of capturing and collection of real-time field information in the modern manufacturing, an affordable approach is introduced [150], which deploys wireless manufacturing technologies (such as RFID) for managing work-in-progress (WIP) inventories in manufacturing job shops.
VI-D Digital twin connects cyber-physical spaces
The digital twin attracts much attention in recent years, some great progresses are achieved. This paper reviews these contributions in terms of accurately modeling and interoperability.
In terms of accurately modeling in manufacturing, current works mainly focus on single-dimension modeling, including structure modeling, attribute modeling, behavior modeling, and process modeling. One finite element analysis approach orienting industrial rail surface wear is constructed [19]. The approach realizes fault monitoring of targets and the iteration of the finite element model by updating the key parameters. A digital twin model for atmospheric and vacuum petrochemical process is proposed [153], in order to achieve safety management and control of the petrochemical processes. For the assembly quality management in the manufacturing process, the surfaces of the assembly parts are modeled, and every stages of assembly process is tracked to reduce errors through online monitoring surface condition in the assembly process [20]. Driven by the increasingly complex and uncertain scheduling process of smart manufacturing system, the work [154] firstly introduces the digital twin modeling of the scheduling processes, and then illustrates some scheduling scheme generated by the algorithm by the twins. In this way, the effectiveness of production line scheduling is enhanced. However, researches of system-level (mixed/multiple-dimensions) modeling are lacking.
Data-driven modeling, knowledge-based modeling, and hybrid modeling are the three most popular modeling methods for the digital twin. A neural network based modeling approach is proposed [155] for the uncertainty and complexity of the digital twin system. The digital twin technique is used into monitoring task of the tool Wear Monitoring of Micro-Milling [156]. The model is based on the existing electrical and mechanical formulas of the three core modules of the micro-milling machine tool. The work [157] presents a digital twin model that makes use of data collected by sensors, so that the performance (accuracy and reliability) can be improved by the knowledge and data analysis. A digital-twin-assisted fault diagnosis using deep transfer learning to analyze the operational conditions of machining tools is proposed[158]. This proposed system has developed an intelligent tool-holder that integrates a k-type thermocouple and cloud data acquisition system over the WiFi module, and the tool-holder can provide better accuracy to demonstrate the performance improvement of milling and drilling operations of cutting tools.
Interoperability supports mutual execution between physical and digital spaces. It is an increasingly attractive topic. However, related researches are still at an early stage, and many problems remain about mutual intelligibility, interoperability, dynamic updating, data heterogeneity, submodel heterogeneity, etc. Currently, some works are achieved that focus more on solving heterogeneous data related problems. The work [159] studies the multi-scale evolution mechanism of product quality for the fusion and use of multi-source heterogeneous data. In the process, heterogeneous data is first classified and reorganized according to attributes and categories. Then according to the data level, it is divided into three categories: macro, medium, and micro. Finally, with the help of the dynamic knowledge base, the concepts of data mapping and rule mapping are used to form a scale dynamic quality map to describe the product quality. Jiang et al.[160] designed a data interaction mechanism that integrates message middleware, memory database and relational database for data interaction in virtual and physical space. The messaging middleware, eg Kafka and MQTT, is used to send production instructions and feedback execution results. The memory database, eg Redis, is used to store real-time operating data during production. The relational databases, eg MySQL, Oracle, SQL Server, and so on, are used for data management for simulation systems.
VI-E Deep learning boosts intelligence
Deep learning methods use computational models that are composed of multiple processing layers to learn representations of data with multiple levels of abstraction [161]. It does not depend on prior data processing and automatically extracts features by using the backpropagation algorithm [162]. In recent years, deep convolutional nets have brought great breakthroughs in the domains of images, video, speech, and games, whereas recurrent neural networks do well in processing sequential data such as text and speech. Moreover, deep learning can speed up the development of IIoT intelligence.
Here deep learning methods are divided into three categories, supervised learning, unsupervised learning and deep reinforcement learning. Supervised learning infers a function from labeled training dataset where each training example is a pair that consists of an input object and a desired output value. Unsupervised learning must self-discover patterns from the training set in which any pre-assigned labels or scores are not provided. Reinforcement learning is an effective method that helps make a sequence of decisions to maximize the rewards by trial and error. This paper demonstrates the enabling effect of deep learning by researches in industrial machine vision as examples.
VI-E1 Supervised learning
Given a certain amount of annotated images, the learned machine vision models are flexible with various types of output, including discrete labels, continuous values for every single pixel or the whole image. The supervised learning provides efficient solutions for parts of the field problems. As an important step for planar measurement, the basic edge detection is achieved with Convolutional Neural Network (CNN) by classifying the pixel [163][164]. The quality of the detected edges are significantly improved in comparison with the traditional edge detectors, e.g., Canny. Compared to traditional fringe projection profilometry with several images, deep neural networks can map a single fringe image to its height/depth image [165][166]. It thus enables accurate 3D surface measurement for moving objects. Measuring the deformation field between two images also benefits the manufacturing processing, both digital image correlation (DIC) and particle image velocimetry (PIV) are modified with a modern deep regression neural network [167, 168, 169]. The neural networks demonstrate improved accuracy and robustness. These geometrical measurement techniques are the foundation for manufacturing, and has been significantly improved with the supervised learning.
VI-E2 Unsupervised learning
One shortage of supervised learning is the need of annotation, which is often expensive. On the contrary, the raw data without labels is quite cheap for most of the industry scenes. As a result, the unsupervised learning based machine vision becomes popular in industry. As a representative of intelligent industrial tasks, defect detection is reviewed with recent progress in this part. One strategy of defect detection is to reconstruct the background image (defect-free), thus the difference between the images and corresponding reconstructed reference implies the defect. A denoise autoencoder networks is used to rebuild the defect free image, thus the reconstruction residuals of the training patches are used as the indicator for defect prediction [170]. Besides, a variational autoencoder (VAE) is also an optional to rebuild the reference images [171]. Similarly, an modified deep convolutional generative adversarial network (DCGAN) is used to reconstruct a given query image such that no defects but only normal textures will be preserved in the reconstruction [172]. The two-stage frameworks build a generative one-class classifier on learned representations which is learned in a self-supervised manner [173][174]. Generally speaking, the supervised methods perform significantly better than the unsupervised methods if sufficient annotation is available [175]. The hybrid two-stage framework takes advantage of both unsupervised representation learning and supervised classification.
VI-E3 Deep reinforcement learning
Intelligent equipment needs the feedback provided by machine vision to automatically make decisions in complex tasks, including robotic assembly, intelligent welding, etc. The challenge of traditional machine vision is to find proper representation which can accurately describe the current state of environment and equipment, such as the robotic assembly with deformable objects [176]. The traditional model-based methods depict the behavior of deformable objects with simplified application-dependent models [177]. Recently, reinforcement learning from high-dimensional sensory inputs is becoming feasible and popular [178][179]. The general deformable objects manipulation might be achieved with high-dimensional reinforcement learning [180]. However, the performance of image-based reinforcement learning (Dreamer [178], PlaNet [181], etc) is far below the optimal performance of the state representation [180]. It points out that learning the dynamics of deformable objects is difficult due to the complex visual observations and dynamics. Despite the challenges of real world [182], the industrial applications of deep reinforcement learning are broaden with the increasing reported cases [183][184].
VI-F Smart hardware builds intelligence foundation
In recent years, smart hardware has realized great breakthroughs in smart sensors and powerful computing devices.
Regarding the sensor technology, the complementary metal-oxide-semiconductor (CMOS) sensors have achieved commercial realization with satisfactory performance efficiency, replacing the charge-coupled devices (CCDs). Due to about 95% quantum efficiency, the latest CMOS sensors become nearly perfect detectors [185]. Event-based camera is a silicon retina which outputs asynchronous spikes with a number of benefits, including extremely low latency, high dynamic range, and very low power consumption [186][187]. For extremely low light imaging conditions, the quanta image sensor (QIS)—a single-photon detector—is worth noting for the unstructured industrial conditions [188][189]. Besides the optical sensors, computational imaging [190] is a wonderful breakthrough in industrial image process. It attempts to balance the processing capabilities of optics and electronics through concurrent design and joint optimization of all elements. It can extend the capability of conventional means, or reduce the measurement cost in comparison to conventional means. Specifically, the computational imaging contains a wide range of techniques, including the phase imaging, computed tomography (CT), magnetic resonance imaging (MRI), time-of-flight (ToF), light-field imaging, time-gated imaging, synthetic aperture radar, coded-aperture imaging, super-resolution, etc. These computational sensors have been adopted widely in industry [191, 192, 193, 194].
The computing capability keeps rapidly increasing due to Moore’s Law and other computing technologies (e.g. domain specific architectures) [195][196]. Therefore, the long data processing time required by computational imaging can be significantly shorter [197][198]. Besides, the platforms equipped with Intel CPU, Nvidia GPU, or Google TPU enable fast training of deep learning models [199][200].
VI-G Cloud/Edge computing inspires novel business and deployment
Cloud computing delivers computing services over the internet to users, and allows users to access and use resources such as computing power, storage, and software applications without the need for owning or managing physical infrastructure. In a cloud computing environment, services are provided on-demand, and users can scale their resources dynamically based on their requirements.
Edge computing is a distributed computing paradigm designed to process data in close proximity to the source of generation, minimizing latency and enhancing the efficiency of applications. Unlike traditional centralized cloud computing, edge computing involves local processing on devices or nearby servers. This proximity allows for real-time or near-real-time data processing, making it particularly suitable for applications such as Internet of Things (IoT) devices and augmented reality, where low latency is crucial.
There are many surveys discussing the technical models of cloud computing [201, 202, 203, 204] and edge computing [205, 206, 207]. This paper will not repeat it, but only focus on the contributions of cloud computing and edge computing in IIoT intelligence, that is, inspiring novel manufacturing business and deployment.
VI-G1 Cloud manufacturing
Cloud manufacturing [33, 34, 35] is a paradigm that leverages cloud computing technologies to enhance and optimize various aspects of the manufacturing process. It represents a shift from traditional, localized manufacturing models to a more connected and dynamic approach. In cloud manufacturing [208], resources, information, and capabilities are provided as services over the internet, enabling a more flexible, scalable, and collaborative manufacturing environment.
Cloud manufacturing as a novel business mode can offer many advantages. Firstly, manufacturing resources, such as equipment, software tools, and computational power, are shared and accessed through the cloud [209, 210]. Advanced manufacturing resource virtualization and sharing strategies play a crucial role in optimizing production processes, improving efficiency, and making informed decisions [211, 212, 213]. This allows for more efficient utilization of resources and reduces the need for individual companies to invest heavily in specialized infrastructure. Secondly, manufacturers can access manufacturing services on a pay-as-you-go basis. This on-demand model allows companies to scale their production up or down based on market demand, without the need for significant capital investment. Thirdly, the cloud enables the collection and analysis of large volumes of data generated throughout the manufacturing process. Data analysis tools can be applied to gain insights into production efficiency, quality control, and overall performance. Finally, cloud manufacturing extends its impact beyond individual factories by integrating the entire supply chain. This integration improves communication, coordination, and responsiveness among suppliers, manufacturers, and distributors.
VI-G2 Cloud-based deployment
Cloud-based deployment refers to the practice of hosting and delivering software applications, services, and resources over the internet through cloud computing infrastructure. In a cloud-based deployment model [214, 215], managers and clients of enterprises access and interact with applications and data remotely, typically through web browsers or client applications, rather than relying on on-premises hardware and infrastructure.
Cloud-based deployment offers a wide range of benefits across various industries. Firstly, cloud computing eliminates the need for significant upfront investments in physical hardware and infrastructure. Enterprises can leverage a pay-as-you-go model, paying only for the computing resources they use. This results in cost savings, especially for small and medium-sized businesses that may not have the capital for extensive IT infrastructure. Secondly, cloud-based deployment allows manufacturers to scale their computing resources based on demand. This is particularly beneficial in smart manufacturing where data volumes can vary, and the need for computational power may fluctuate. Scalability ensures that resources are readily available to handle increased workloads during peak production times.. Thirdly, cloud-based deployment enable remote monitoring and management of manufacturing processes. This is particularly relevant in smart manufacturing, where sensors and IoT devices generate real-time data. Remote access allows stakeholders to monitor production, identify issues, and make informed decisions from anywhere with an internet connection. Finally, cloud computing provides an environment for manufacturing enterprises to experiment and adopt innovative technologies. This includes integrating artificial intelligence, machine learning, and other emerging technologies into smart manufacturing processes. Cloud-based platforms offer the computational power needed for training and deploying advanced models.
VI-G3 Edge-based deployment
Edge-based deployment refers to the strategic placement of computing resources, data processing, and analysis capabilities closer to the source of data generation, typically at the network edge. Unlike cloud-based deployment where industrial data is sent to centralized servers for processing, edge-based deployment shifts some of the processing tasks to local devices or edge servers. This approach is particularly beneficial in industrial settings where real-time processing, low latency, and reliable connectivity are crucial for operational efficiency and decision-making [216, 217, 218].
Compared with cloud-based deployment, edge-based deployment can obtain advantages in low latency, bandwidth optimization, autonomous operation. Firstly, edge-based deployment processes data closer to the source reduces latency, enabling real-time or near-real-time responses. Secondly, edge-based deployment reduces the need to transmit large volumes of raw data to the cloud, optimizing bandwidth usage. This is especially important in scenarios where network bandwidth is limited or expensive. Thirdly, edge-Based deployment can operate autonomously even without continuous connectivity to the cloud. This ensures that critical functions can continue in the event of network disruptions, enhancing system reliability.
VII Open Challenges and Future Trends
To identify existing challenges and circumvent potential misleading directions of IIoT intelligence, we think about questions:
  • What kinds of capabilities are required by future manufacturing?
  • Which capabilities have been established?
  • How to establish or improve the capabilities that are required by future manufacturing?
After careful argumentation, We obtain keywords digital control, deterministic response, cost-friendly operation and deployment, IIoT intelligence proliferation to answer the first question. The current IIoT intelligence has achieved certain capabilities, however, there is still room for improving them.
VII-A Digital Control
Digital control refers to the use of digital technologies and advanced control systems to monitor, manage and optimize various aspects of the manufacturing processes. This approach leverages digital information, connectivity, and data analysis to enhance control and decision-making in real-time.
Current IIoT intelligence can support digital control to a certain extent. In the current architecture of IIoT intelligence, industrial sensors are integrated to provide real-time information about the production processes, and the data collected by sensors are processed for analysis. The automation systems in software layer, which are powered by programmable logic controllers (PLCs) and distributed control systems (DCS), allow for precise control over manufacturing equipment and processes based on the real-time data analysis. Remote control and monitoring also facilitate digital control by adjust manufacturing processes from remote locations.
However, the current level of digital control is still far from achieving its ideal vision of providing real-time visibility into all aspects of the manufacturing process, and any remote or local commands can be quickly executed in visual virtual spaces and physical manufacturing processes. The real-time visibility indicates that future manufacturing has the ability to monitor the status of equipment, track the flow of materials and analyze production data as it is generated. The immediate response for commands indicates that the effective interaction across multiple layers of IIoT intelligence is realized, as well as the cooperation among robots, sensors, computation resources and network resources. More importantly, once the ideal vision of digital control is realized, managers can grasp various aspects of the manufacturing environment through real-time monitoring, adjustment, optimization, and control.
It is challenging to realize the vision, but the digital twins connecting cyberspace and physical space can provide some opportunities after overcoming some difficulties. Firstly, due to the frequent interaction between digital twin models and industrial software to exchange real-time industrial data/commands, it is necessary to address interoperability issues between software and models, as well as standardization issues between communication protocols in industrial environments. Secondly, industrial datasets often suffer from issues such as noise, outliers, and missing values, impacting data quality, inconsistent data formats and units across different sources. Poor data quality leads to inaccurate analyses and unrealiable conclusions. Therefore, it is non-trivial to generate high-quality industrial data based on the practical data, in order to enable intelligent data analysis. Thirdly, at the beginning of digital twin design, due to our inability to collect/generate a perfect dataset that includes the behavior of modeled objects under all conditions, the performance of digital twins will decrease with the increase of invisible conditions. Therefore, it is necessary to design update mechanisms for digital twins to maintain the high-performance operation of the digital twin system.
VII-B Deterministic response
The capability of deterministic response signifies an ability to produce predictable and certain outcomes or behavior based on specific inputs or initial conditions. With deterministic response, there is no randomness or uncertainty, and the future state of the system can be precisely determined based on its current state and the inputs it receives.
In the current IIoT intelligent architecture, Time Sensitive Networking (TSN) ensures deterministic response for network traffics. However, as the industrial digital transformation process accelerates, many time-critical computing tasks are generated by networked devices, such as sensor data analysis and motion planning for robots. Timely completion of the computing tasks is crucial for ensuring the correct and efficient operations of manufacturing systems. Moreover, as the cooperative relationships between robots, sensors, software and models in IIoT become more complex, there is an increasing expectation for the response delays of certain time-critical computing tasks to be deterministic. Current end-edge-cloud cooperative computing architecture, which relies on best-effort networks to forward data, leads to uncertainty and unpredictability in the response time of IIoT computing tasks. Therefore, the issue of ensuring the deterministic response of IIoT computing tasks has not been studied yet. It will become a future research direction.
VII-C Cost-friendly operation and deployment
Cost-friendly operation refers to the practice of managing and conducting business operations in a manner that optimizes costs, enhances efficiency, and ensures financial sustainability. Cloud manufacturing has provided such a cost-friendly operation that allows on-demand manufacturing. However, most researchers focus on fine-grained manufacturing resource scheduling in the perspective of cloud platform instead of manufacturing resource providers. Therefore, attention should be given to the pricing and supply strategies of manufacturing resources for manufacturing enterprises, in order to enable them to obtain higher profits at lower operating costs.
Cost-friendly deployment refers to the strategic and efficient implementation of systems or technologies with a focus on optimizing costs while achieving desired outcomes. The cost-friendly deployment of manufacturing processes is often realized by professionals who specialize in various aspects of industrial engineering, production management, and process optimization. These experts contribute their knowledge and skills to enhance efficiency, reduce costs, improve product quality, and streamline overall operations. We won’t interfere with the deployment optimization of manufacturing processes.
As the digital development of smart manufacturing, cost-friendly deployment of computation and storage resources is also required to ensure industrial data storage and analysis. Cloud-based deployment and edge-based deployment are able to provide cost-friendly deployment strategies for industrial big data and systems. However, based on the analysis in digital control and deterministic response, we can know that there are many time-critical tasks in IIoT. Both cloud-based deployment and edge-based deployment will introduce unexpected and unpredictable latency due to the usage of best-effort networks. Therefore, a novel local deployment strategy for computation and storage resources should be studied in the future, as well as the resource scheduling mechanisms, in order to ensure the deployment and scheduling strategies can satisfy latency requirement of all IIoT tasks within a constrained budget.
VII-D IIoT intelligence proliferation
The intelligence proliferation refers to the development or creation of new intelligent systems, models, and algorithms through AI, to proceed IIoT tasks or emergency situation. The reasons why the intelligent proliferation is required in future smart manufacturing include:
  • Industrial manufacturing involves intricate production processes, equipment, and systems, making it challenging to accurately describe and address all possible scenarios through traditional manual rules. The diversity and complexity of smart manufacturing environments require systems to adapt to different conditions and requirements.
  • Designing and optimizing specialized algorithms for each possible scenario is a time-consuming task. Relying solely on manually designed case-by-case algorithms may not meet the demands of rapid changes and complexity in dynamic manufacturing environments.
  • The potential scenarios for industrial manufacturing are diverse, and some situations may be unpredictable. Traditional manual design methods are difficult to fully cover all possible scenarios.
  • Intelligent proliferation can learn and understand patterns and regularities in industrial environments through accumulating industrial knowledge. When encountering emergency situations, intelligent proliferation system can automatically generate corresponding solutions to reduce the harm caused to the production process.
  • Intelligent emergence allows systems to automatically generate adaptive processing algorithms when facing new types of data tasks, significantly reducing the workload of independent development for each scenario. This adaptability can greatly enhance the system’s flexibility and efficiency.
  • Intelligent emergence systems possess adaptability and learning capabilities, continuously optimizing their models and algorithms based on real-time industrial data. This enables the system to continually adapt to changing industrial environments and production requirements.
Inspired by the development of ChatGPT, which can understand and generate human-like text in conversations, retain context for coherent responses, perform tasks like answering questions and creative writing, and provide information based on its knowledge, a large industrial generative intelligent model has potential to achieve IIoT intelligence proliferation, and will be one of the future research directions.
VIII Conclusion
Due to the fiercely competitive business environment, increasingly personalized customization needs, and rising labor costs, the manufacturing industry is eager to undergo digital transformation. At the same time, many data-driven and knowledge-based intelligent technologies are emerging, lighting the way for intelligent manufacturing. IIoT intelligence refers to a series of techniques, methods, productions and platforms that can be adopted throughout the entire value chain to establish capabilities of digital connection and perception, intelligent analysis and cognition, immediate decisions, etc. It can provide innovative and efficient solutions for all aspects of the manufacturing value chain. Therefore, IIoT intelligence will drive the transformation and upgrading of smart manufacturing in terms of functionality, operation, deployment, and application.
We have designed a hierarchical development architecture for IIoT intelligence to build a systematic understanding of how IIoT intelligence deals with the issues in the transformation process of smart manufacturing. The architecture consists of equipment layer, networking layer, software layer, modeling layer, and analysis and optimization layer. The equipment layer builds the foundation of IIoT intelligence. The networking layer realizes ubiquitous connection among human-machine-thing. The software layer provides digital expression of industrial processes. The modeling layer achieves accurately modeling and interoperability improvement. The analysis and optimization layer supports case-by-case algorithm design to analyze industrial data and optimize industrial processes. In addition, we also closely observe the lighthouse factory to demonstrate the positive effects of technologies at each layer.
Following that, we identify seven kinds of technologies that accelerate the development of IIoT intelligence, including industrial robots, machine vision, networking technologies, digital twins, deep learning, smart hardware, cloud/edge computing. Industrial robots can become intelligent labors to overtake dirty and dangerous tasks. Machine vision has potential to be the eyes of industry. Networking facilitates ubiquitous connectivity. Digital twins connect cyberspace and physical spaces. Smart hardware builds the foundation of intelligence. Cloud/Edge computing inspires novel business and deployment.
Finally, we explore the open challenges and discuss the future research directions from four aspects, including developing digital twins to realize ideal vision of digital control, exploring deterministic response strategies and mechanisms for IIoT computing tasks, optimizing computation and network resources deployment to reduce costs, and researching industrial generative AI large model to realize intelligence proliferation.